Part C

The Expert TA I Human-like Grading, Automated Home Student: amele7@mea Work and Energy Begin Date: 1/19/2019 1200:00 AM-Due Date: 1/24/2019 11-59.00 PM End Date: 2/14/2019 11:00.00 PM (7%) Problem 13: of 26 N. A 85-kg cosHountry skier is climbinga 3.15, slope at a constant speed of 2.s mind encounten air resia ce İ ψ 33% Part (a) Fnd his power output, in watts, for he work be des against the gravitational force and air resistance. P-154 Correct ed| |-V33% Part (b) what is the magnitude ofthe average force, in newtons, des he exert backward on the snow t accomplish this? F-718 Correct ted > st 33% Part (c) r he continues to exert this force and to experience te same air resistance when he reaches a level area, how kng will it take him reach a vekocity of 5.5 ms? Grade Summary Podential eted ctod coranasin acos atano acotan0 sinh0 Anemps maining:12 (1% per athempt) MacBook Air 20 s 808 1 5 0
Answer #1

F = applied force = 71.8 N

f = force by air resistance = 26 N

m = mass of the skier = 85 kg

a = acceleration of the skier = ?

force equation for the motion of the skier is given as

F - f = ma

71.8 - 26 = 85 a

a = 0.54 m/s2

vi = initial velocity = 2.15 m/s

vf = final velocity = 5.5 m/s

t = time of travel

using the equation

vf = vi + a t

5.5 = 2.15 + (0.54) t

t = 6.2 sec